# The Art of Saying "Maybe": A Conformal Lens for Uncertainty Benchmarking in VLMs

### v2

@@ 6.5 Domain-Specific Uncertainty Performance
VLMs show better uncertainty calibration on datasets where visual elements complement rather than dominate reasoning. ScienceQA, with images reinforcing textual concepts, yields high accuracy (75.2% average) with well-calibrated uncertainty (2.1 average set size). Conversely MathVision, requiring precise extraction of visual quantitative information - proves challenging for uncertainty calibration (4.5 average set size despite lower accuracy).
Figure 8: Correlation matrix between model size, accuracy, set size, and coverage rate. Diagonals represent distribution for values of each metric.
In specialized domains like medical visual reasoning (WorldMedQAV), uncertainty calibration remains stable within parameter brackets despite varying accuracy levels. This suggests domain-specific visual expertise and uncertainty awareness develop somewhat independently; models may recognize domain-specific features without being well-calibrated about their confidence, or vice versa.
@@ 6.6 Entropy-Based Uncertainty Analysis
Table 4 reports entropy-based uncertainty scores across models and datasets. The results closely align with our conformal prediction findings: larger models generally achieve lower entropy (better calibration), with the same Gemma 12B vs. 4B anomaly observed in set size metrics. The Gemma family overall ranks strongest, with all variants appearing among the top performers. Domain-level trends are also consistent - ScienceQA yields the lowest entropy values, while MathVision remains most challenging. These results validate our primary analysis and confirm that our conclusions hold across different uncertainty estimation approaches.
@@ 7 Conclusion
This work presents a comprehensive conformal uncertainty benchmarking study for Vision-Language Models, revealing important correlations between model scale, accuracy, and uncertainty quantification capabilities. Our findings demonstrate that larger models not only achieve higher accuracy but also produce better-calibrated uncertainty estimates; they know better what they don’t know.
@@
While this study provides valuable insights into uncertainty quantification for VLMs, several limitations should be acknowledged. First, our benchmarking was restricted to multiple-choice datasets due to the computational constraints of conformal prediction methods, which require well-defined prediction sets. Future work should explore extending these techniques to generative tasks where the output space is unbounded, potentially through methods that quantify uncertainty in free-form text generation. Second, our proprietary model analysis was limited to GPT-based models, as other leading systems like Claude and Gemini do not currently expose token-level log probabilities necessary for our conformal scoring functions. As these capabilities become available in more proprietary models, the benchmarking framework could be readily extended to provide a more comprehensive landscape of uncertainty quantification across the industry.
@@ Appendix A Prompt Design Specifications
This section documents the complete prompt engineering framework used in our experiments. The two-layer prompting strategy consists of:
  • System messages that establish the model’s role and response constraints
  • Zero-shot instructions that provide task-specific guidance with MCQ questions and options, while maintaining output consistency
The combination ensures standardized evaluation conditions across all datasets while preserving each task’s unique requirements. Tables 5 and 6 detail the exact formulations. All datasets used standardized zero-shot instructions following the pattern mentioned in Table 6.
Other datasets used identical phrasing, adjusted only for:
  • Domain specificity (e.g., “scientific diagram” for AI2D, “medical image” for WorldMedQAV)
  • Option letter range (A-F for most, up to A-J for MMMU-Pro)
@@
Table 5: System Prompts for each dataset in the VLM evaluation
Dataset Example Prompt
Please select the correct answer from the given options.
Only respond with the option letter (A, B, C, D, E).
{QUESTION}
{OPTIONS}
Table 6: Representative zero-shot prompt example
Appendix B Dataset Statistics and Preprocessing
This section documents key dataset characteristics and preprocessing steps taken to ensure evaluation consistency across all benchmarks.
B.1 Option Distribution Normalization
We standardized answer options across all datasets to maintain consistency in evaluation:
  • Added "I don’t know" and "None of them" options where needed to ensure uniform option counts within each dataset
  • Maintained original option ordering (A, B, C,…) without randomization
B.2 Multimodal Sample Handling
To ensure fair evaluation across all models:
  • Excluded questions with multiple input images (4.9% of total samples from MMMU dataset) since not all evaluated VLMs support this feature
  • Verified all remaining samples contain exactly one image-question pair
B.2.1 Option Balance Analysis
The option distributions reveal distinct patterns across datasets:
  • Balanced Datasets: AI2D, MathVision, WorldMedQAV, and MMMU-Pro show approximately uniform answer distributions (Figures 9(a), 9(c), 9(d), 9(f))
  • Skewed Datasets: ScienceQA and MMMU show a higher frequency in earlier options(Figures: 9(b), 9(e)).
Table 7: Final test set statistics after preprocessing
Dataset Description Samples Options
@@ B.2.1 Option Balance Analysis
Figure 9: Answer option (ground truth) distributions across all six benchmark datasets. Each subplot shows the frequency of correct answers.

### v3

@@ 6.5 Domain-Specific Uncertainty Performance
VLMs calibrate uncertainty better when visuals support reasoning, not lead it. ScienceQA, where images back text, scores high accuracy (75.2% average) and tight sets (2.1 average). MathVision, needing exact visual number pulls, challenges calibration (4.5 average set size, lower accuracy).
In medical tasks (WorldMedQAV), calibration stays steady within size groups despite accuracy swings. This implies domain visual skills and uncertainty awareness evolve separately, models might spot features without gauging confidence accurately, or the reverse.
@@ 6.6 Entropy-Based Uncertainty Analysis
Table 3 shows entropy scores, mirroring conformal results: larger models demonstrate lower entropy (better calibration), with the Gemma 12B vs. 4B outlier in set sizes. Gemma leads overall, with all sizes in top spots. Domains follow suit; ScienceQA exhibit the lowest entropy, and MathVision the highest. This backs our main findings and extends them to alternative uncertainty methods.
Table 4: Set Size ($\downarrow$) for closed-source VLMs using the instruction-guided likelihood proxy.
Task Characteristic Datasets Scoring Adaptation Rationale
Low Ambiguity / Fact-Based ScienceQA, AI2D LAC Probability mass is concentrated; LAC yields the smallest sets while maintaining 90% coverage, reducing unnecessary options.
High Ambiguity / Complex Reasoning MathVision, MMMU-Pro APS Distributions are diffuse; APS accumulates probability mass adaptively, improving robustness under reasoning uncertainty.
Differential / Binary Ambiguity WorldMedQAV MS Tasks often involve two leading hypotheses; MS models confidence gaps, avoiding low-relevance tail options.
Table 5: Task-adaptive scoring function recommendations based on task ambiguity and uncertainty structure.
6.7 Instruction-Guided Uncertainty Proxies
Table 6: Alignment between GPT-4o-mini’s true logprobs and instruction-guided proxy likelihoods across six benchmarks. Moderate-to-strong correlation validates proxy-based uncertainty for closed-source VLMs.
To evaluate closed-source VLMs without logprob access, we adopt an instruction-guided likelihood proxy where models output option-wise numerical likelihoods in a strict JSON format, used as an uncertainty signal.
We validate this proxy on GPT-4o-mini by comparing it against true token-level logprobs. Table 6 shows moderate-to-strong alignment (Pearson up to 0.47, Spearman up to 0.71, $\sim$74% directional agreement), indicating that the proxy preserves relative uncertainty structure.
Using this proxy, we evaluate GPT-4o-mini, Gemini-2.5-Flash, and Claude Haiku 4.5. Set-size results in Table 4 mirror earlier findings: stronger models yield smaller, better-calibrated sets, while visually intensive datasets (MathVision, MMMU-Pro) produce larger sets for weaker models. Accuracy and coverage trends remain unchanged (Appendix Tables 11, 12).
6.8 Task-Adaptive Scoring Functions
We study how conformal scoring functions vary with task characteristics and derive lightweight selection guidelines. Using prediction entropy as a proxy for task ambiguity (Table 3), we link ambiguity with set efficiency and coverage behavior (Tables 213).
Table 5 summarizes consistent patterns: LAC performs best in low-ambiguity, fact-based tasks; APS is more robust in high-ambiguity reasoning domains; and MS is effective for differential or near-binary uncertainty. These results suggest that task-aware scoring selection yields better-calibrated uncertainty than a uniform choice.
@@ 7 Conclusion
This work presents a comprehensive conformal uncertainty benchmarking framework for Vision-Language Models (VLMs), systematically analyzing how uncertainty behaves across model scales, architectures, and domains. Using conformal prediction, a distribution-free, model-agnostic method with formal guarantees, we evaluated 18 state-of-the-art VLMs over six multimodal benchmarks and three scoring functions.
Our results reveal clear trends: larger models not only achieve higher accuracy but also produce more calibrated uncertainty estimates; they know better what they do not know. Scaling enhances both performance and confidence reliability across model families such as Qwen-VL, Gemma, and InternVL.
Overall, this study establishes a principled foundation for multimodal uncertainty evaluation and highlights conformal prediction as a practical, theoretically grounded tool for measuring trust in large VLMs. By quantifying when models are unsure, our approach moves toward building safer, more accountable, and trustworthy AI systems, enabling future research on dynamic calibration and adaptive uncertainty in real-world applications.
@@
This work offers key insights on VLM uncertainty, but limitations persist. First, we stuck to multiple-choice datasets, as conformal methods need defined prediction sets amid compute limits. Future efforts could adapt for generative tasks with unbounded outputs, like uncertainty in open text via specialized techniques.
Second, while we expanded our proprietary model evaluation beyond GPT to include Gemini-2.5-Flash and Claude Haiku 4.5, this extension required an instruction-guided likelihood proxy rather than direct token-level logprobs. Most commercial VLM providers - including Google, Anthropic, and others - do not expose token-level log probabilities through their APIs, a design choice that fundamentally limits principled uncertainty quantification. Our proxy method, which prompts models to self-report option-wise likelihoods in a structured format, demonstrates moderate-to-strong alignment with true logprobs (Section 6.7) but remains an indirect approximation. This proxy inevitably introduces noise from the model’s self-assessment biases, instruction-following variability, and the inherent gap between generated likelihood estimates and actual next-token distributions. If providers were to expose token-level logprobs - as GPT models partially do - our conformal framework could be applied directly, yielding more precise and theoretically grounded uncertainty quantification. Until such interfaces become standard, uncertainty benchmarking for closed-source VLMs will rely on proxy methods that, while practically useful, lack the formal guarantees of direct logprob-based conformal prediction.
@@ Appendix A Prompt Design Specifications
To ensure consistent and reproducible evaluation across all datasets and models, we employed a standardized two-layer prompting strategy. This approach separates domain-specific role assignment (via system messages) from task-oriented instructions (via zero-shot prompts), minimizing variability while adapting to each benchmark’s unique requirements. System messages define the model’s persona and output constraints, promoting focused responses. Zero-shot instructions provide contextual framing without examples, encouraging natural reasoning. This design facilitates reliable extraction of token-level probabilities for conformal prediction, as all responses are constrained to single-letter outputs (e.g., A-E). Below, we detail the formulations used, which were applied uniformly except for minor adjustments in option ranges (e.g., up to J for MMMU-Pro).
Tables 8 and 9 present the exact system messages and a representative zero-shot example, respectively. These prompts were tested for compatibility across all 18 VLMs, ensuring no leakage of solving strategies that could artificially inflate confidence estimates.
Appendix B Dataset Statistics and Preprocessing
Dataset Description Samples Options
Table 7: Final test set statistics after preprocessing.
Our evaluation relies on six diverse multimodal datasets, each preprocessed to ensure uniformity and compatibility with the conformal prediction framework. Preprocessing focused on standardizing option distributions, handling multimodal inputs, and balancing calibration/test splits (50/50 per dataset). We excluded multi-image samples (e.g., 4.9% from MMMU) to accommodate model limitations, resulting in single-image-question pairs for all instances. Additionally, we normalized options by adding neutral choices (e.g., “I don’t know”) where needed, without altering ground-truth labels or randomization to preserve natural distributions. These steps mitigate biases in uncertainty estimates, such as skewed option frequencies that could inflate set sizes under conformal scoring.
Table 7 summarizes the final test set sizes and option ranges post-preprocessing. Figure 7 visualizes ground-truth answer distributions, revealing patterns like balance in AI2D/MathVision (uniform across options) versus skew in ScienceQA/MMMU (favoring early letters A-C). Such insights highlight domain-specific challenges: balanced datasets like MMMU-Pro test broad expertise, while skewed ones may reflect annotation biases, influencing conformal coverage in reasoning-heavy tasks.
Appendix C Additional Results and Analyses
Figure 6: Set size distribution across different uncertainty methods. Distributions highlight LAC’s compactness and APS’s sensitivity to domain ambiguity.
Dataset Example Prompt
Please select the correct answer from the given options.
Only respond with the option letter (A, B, C, D, E).
{QUESTION}
{OPTIONS}
Table 8: Representative zero-shot prompt example. Similar phrasing was adapted for other datasets, with adjustments for domain (e.g., “math problem” for MathVision) and option count.
@@ Appendix C Additional Results and Analyses
Table 9: System Prompts for each dataset in the VLM evaluation.
@@
Figure 7: Answer option (ground truth) distributions across all six benchmark datasets. Each subplot shows the frequency of correct answers, illustrating balance (e.g., AI2D) versus skew (e.g., ScienceQA).
This appendix extends the main results with supplementary metrics and visualizations, providing deeper validation of our conformal prediction framework and uncertainty patterns. We focus on coverage guarantees (essential for statistical reliability), set size distributions (to assess scoring function variability), and inter-metric correlations (to reinforce scaling trends). These analyses confirm the robustness of our findings: larger models exhibit tighter uncertainty bounds, while domain complexity (e.g., MathVision) amplifies set sizes across methods.
C.1 Coverage Rate Tables and Plots
Coverage rate (CR) verifies the conformal framework’s $1-\alpha=90\%$ guarantee, measuring the proportion of true labels included in prediction sets. As expected, CR remains stable and above threshold for most models/datasets, with minor dips in high-complexity tasks like MMMU-Pro (due to diffuse probabilities). This supports the method’s distribution-free applicability to VLMs, even under varying architectures. Table 13 details per-model/dataset CR across scoring functions, while Figure 8 compares distributions, highlighting LAC’s edge in maintaining coverage with compact sets, as depicted in main text Table 2.
C.2 Set Size Distributions by Method
To explore scoring function behaviors beyond averages (main text Figure 2), we analyze set size distributions. These reveal APS’s adaptability in ambiguous cases (broader tails for MathVision) versus LAC’s consistency (narrower peaks). Such variability underscores the need for task-specific scoring selection, aligning with our observation that LAC suits VLM confidence estimation best. Figure 6 supplements the main analysis, showing histogram overlaps that confirm inverse accuracy-set size correlations (as shown in Figure 1).
C.3 Distribution-Shift Robustness
Table 10: Cross-dataset transfer coverage rates and set sizes using APS at $(1-\alpha)=90\%$. Diagonal entries are in-distribution; off-diagonal entries show transfer performance.
Table 11: Coverage Rate ($\uparrow$) for closed-source VLMs using the instruction-guided likelihood proxy.
Figure 8: Coverage rates across datasets and models. All scoring methods maintain at least 90% coverage in most cases, with APS showing slight variability in reasoning tasks.
We assess CP robustness under distribution shift via cross-dataset transfer using APS at $(1-\alpha)=90\%$, calibrating on one dataset and testing on another.
Coverage remains stable under shift: 93.5% in-distribution vs. 92.8% out-of-distribution. Most transfers are conservative (e.g., MMMU-Pro $\rightarrow$ ScienceQA), while shifts from easier to harder datasets (ScienceQA $\rightarrow$ MMMU-Pro) show minimal degradation, indicating strong generalization. Table 10 reports coverage and average set size across all source–target pairs; diagonal entries denote in-distribution results.
C.4 Closed-Source VLM Evaluation
Table 12: Accuracy ($\uparrow$) of closed-source VLMs using the instruction-guided likelihood proxy across six benchmarks.
This section presents supplementary results for closed-source VLMs (GPT-4o-mini, Gemini-2.5-Flash, Claude Haiku 4.5) evaluated using the instruction-guided likelihood proxy method described in Section 6.7 of the main text. Since these proprietary models do not expose token-level logprobs, we instructed models to output option-wise numerical likelihoods in a structured JSON format, which serves as a proxy uncertainty signal for conformal prediction.
Table 12 presents accuracy performance, demonstrating that accuracy trends remain consistent with open-source models across all benchmarks. Table 11 shows coverage rates, verifying that the conformal prediction framework maintains its statistical guarantees even when applied through the proxy method.
C.5 Correlation and Relationships Between Metrics
Finally, we quantify relationships between key metrics to validate scaling effects. The correlation matrix in Figure 9 shows strong negative ties between accuracy and set size ($r\approx-0.75$), positive links to model size, and stable coverage. Diagonals illustrate metric spreads, e.g., set sizes clustering below 3 for >10>10B models. These reinforce main findings: uncertainty calibration improves with scale, offering a "conformal lens" for VLM benchmarking.
Figure 9: Correlation matrix between model size, accuracy, set size, and coverage rate. Negative correlation between accuracy and set size is clearly observed. Diagonals represent distributions for each metric.